# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ Acknowledgements.
We thank J. Venturini, C. Mordasini, Th. Henning, H. Klahr, and E. Pallé for fruitful discussion. We further thank the contributions and discussions of all participants of the ”Density Matters” Ringberg workshop which helped to shape this work. R.B. acknowledges the financial support from DFG under Germany’s Excellence Strategy EXC 2181/1-390900948, Exploratory project EP 8.4 (the Heidelberg STRUCTURES Excellence Cluster) and from the European Research Council under the European Union’s Horizon 2020 research and innovation program under grant agreement No. 832428-Origins. R.L. acknowledges support provided by NASA through the NASA Hubble Fellowship grant #HST-HF2-51559.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. C.D. acknowledges support from the Swiss National Science Foundation under grant TMSGI2_211313. Parts of this work has been carried out within the framework of the NCCR PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. Calculation were performed on the Horus cluster at the University of Bern and on the computation facilities at the Max Planck Institute for Astronomy. The plots shown in this work were generated using matplotlib (Hunter, 2007) and the analysis made use of the scikit-learn (Pedregosa et al., 2011) and scipy (Virtanen et al., 2020) packages.References
- Acuña et al. (2021) Acuña, L., Deleuil, M., Mousis, O., et al. 2021, A&A, 647, A53
- Affolter et al. (2023) Affolter, L., Mordasini, C., Oza, A. V., Kubyshkina, D., & Fossati, L. 2023, A&A, 676, A119
- Aguichine et al. (2021) Aguichine, A., Mousis, O., Deleuil, M., & Marcq, E. 2021, ApJ, 914, 84
- Alibert & Benz (2017) Alibert, Y. & Benz, W. 2017, A&A, 598, L5
- Alibert et al. (2005) Alibert, Y., Mordasini, C., Benz, W., & Winisdoerffer, C. 2005, A&A, 434, 343
- Andrews et al. (2018) Andrews, S. M., Terrell, M., Tripathi, A., et al. 2018, ApJ, 865, 157
- Andrews et al. (2010) Andrews, S. M., Wilner, D. J., Hughes, A. M., Qi, C., & Dullemond, C. P. 2010, ApJ, 723, 1241
- Astudillo-Defru et al. (2020) Astudillo-Defru, N., Cloutier, R., Wang, S. X., et al. 2020, A&A, 636, A58
- Bajgain et al. (2015) Bajgain, S., Ghosh, D. B., & Karki, B. B. 2015, Nature Communications, 6, 8578
- Baraffe et al. (2015) Baraffe, I., Homeier, D., Allard, F., & Chabrier, G. 2015, A&A, 577, A42
- Beatty et al. (2024) Beatty, T. G., Welbanks, L., Schlawin, E., et al. 2024, ApJ, 970, L10
- Benneke et al. (2024) Benneke, B., Roy, P.-A., Coulombe, L.-P., et al. 2024ArXiv e-prints
- Berger et al. (2020) Berger, T. A., Huber, D., Gaidos, E., van Saders, J. L., & Weiss, L. M. 2020, AJ, 160, 108
- Bodenheimer & Pollack (1986) Bodenheimer, P. & Pollack, J. B. 1986, Icarus, 67, 391
- Bonfanti et al. (2024) Bonfanti, A., Brady, M., Wilson, T. G., et al. 2024, A&A, 682, A66
- Bonomo et al. (2023) Bonomo, A. S., Dumusque, X., Massa, A., et al. 2023, A&A, 677, A33
- Borucki et al. (2010) Borucki, W. J., Koch, D., Basri, G., et al. 2010, Science, 327, 977
- Boukrouche et al. (2021) Boukrouche, R., Lichtenberg, T., & Pierrehumbert, R. T. 2021, ApJ, 919, 130
- Bower et al. (2022) Bower, D. J., Hakim, K., Sossi, P. A., & Sanan, P. 2022, The Planetary Science Journal, 3, 93
- Bower et al. (2018) Bower, D. J., Sanan, P., & Wolf, A. S. 2018, Physics of the Earth and Planetary Interiors, 274, 49
- Brinkman et al. (2023) Brinkman, C. L., Weiss, L. M., Dai, F., et al. 2023, AJ, 165, 88
- Burn & Mordasini (2024) Burn, R. & Mordasini, C. 2024, in Handbook of Exoplanets, 2nd edn., ed. H. Deeg, J. A. Belmonte, & R. E. Pudritz (Cham: Springer International Publishing)
- Burn et al. (2024) Burn, R., Mordasini, C., Mishra, L., et al. 2024, Nature Astronomy, 8, 463
- Burn et al. (2021) Burn, R., Schlecker, M., Mordasini, C., et al. 2021, A&A, 656, A72
- Chabrier et al. (2019) Chabrier, G., Mazevet, S., & Soubiran, F. 2019, ApJ, 872, 51
- Chachan & Stevenson (2018) Chachan, Y. & Stevenson, D. J. 2018, ApJ, 854, 21
- Chakrabarty & Mulders (2024) Chakrabarty, A. & Mulders, G. D. 2024, ApJ, 966, 185
- Chambers (1999) Chambers, J. E. 1999, MNRAS, 304, 793
- Charnay et al. (2015) Charnay, B., Meadows, V., & Leconte, J. 2015, ApJ, 813, 15
- Chassefière (1996) Chassefière, E. 1996, Icarus, 124, 537
- Chen et al. (2024) Chen, D.-C., Mordasini, C., Xie, J.-W., Zhou, J.-L., & Emsenhuber, A. 2024, A&A, 687, A25
- Chen & Rogers (2016) Chen, H. & Rogers, L. A. 2016, ApJ, 831, 180
- Cherubim et al. (2024) Cherubim, C., Wordsworth, R., Hu, R., & Shkolnik, E. 2024, ApJ, 967, 139
- Cloutier et al. (2024) Cloutier, R., Greklek-McKeon, M., Wurmser, S., et al. 2024, MNRAS, 527, 5464
- Cloutier & Menou (2020) Cloutier, R. & Menou, K. 2020, AJ, 159, 211
- Coleman et al. (2019) Coleman, G. A. L., Leleu, A., Alibert, Y., & Benz, W. 2019, A&A, 631, A7
- Connolly (2009) Connolly, J. A. D. 2009, Geochemistry, Geophysics, Geosystems, 10, Q10014
- Coppari et al. (2021) Coppari, F., Smith, R. F., Wang, J., et al. 2021, Nature Geoscience, 14, 121
- Cumming et al. (1999) Cumming, A., Marcy, G. W., & Butler, R. P. 1999, ApJ, 526, 890
- Damasso et al. (2023) Damasso, M., Rodrigues, J., Castro-González, A., et al. 2023, A&A, 679, A33
- Dittkrist et al. (2014) Dittkrist, K.-M., Mordasini, C., Klahr, H., Alibert, Y., & Henning, T. 2014, A&A, 567, A121
- Dorn et al. (2015) Dorn, C., Khan, A., Heng, K., et al. 2015, A&A, 577, A83
- Dorn & Lichtenberg (2021) Dorn, C. & Lichtenberg, T. 2021, ApJ, 922, L4
- Dorn et al. (2018) Dorn, C., Mosegaard, K., Grimm, S. L., & Alibert, Y. 2018, ApJ, 865, 20
- Dorn et al. (2017) Dorn, C., Venturini, J., Khan, A., et al. 2017, A&A, 597, A37
- Emsenhuber et al. (2023a) Emsenhuber, A., Burn, R., Weder, J., et al. 2023a, A&A, 673, A78
- Emsenhuber et al. (2023b) Emsenhuber, A., Mordasini, C., & Burn, R. 2023b, The European Physical Journal Plus, 138, 181
- Emsenhuber et al. (2021) Emsenhuber, A., Mordasini, C., Burn, R., et al. 2021, A&A, 656, A69
- Faik et al. (2018) Faik, S., Tauschwitz, A., & Iosilevskiy, I. 2018, Computer Physics Communications, 227, 117
- Fischer et al. (2011) Fischer, R. A., Campbell, A. J., Shofner, G. A., et al. 2011, Earth and Planetary Science Letters, 304, 496
- Fortier et al. (2013) Fortier, A., Alibert, Y., Carron, F., Benz, W., & Dittkrist, K.-M. 2013, A&A, 549, A44
- Fulton & Petigura (2018) Fulton, B. J. & Petigura, E. A. 2018, AJ, 156, 264
- Fulton et al. (2017) Fulton, B. J., Petigura, E. A., Howard, A. W., et al. 2017, AJ, 154, 109
- Gaidos et al. (2024) Gaidos, E., Ali, A., Kraus, A. L., & Rowe, J. F. 2024, MNRAS, 534, 3277
- Giacalone et al. (2022) Giacalone, S., Dressing, C. D., Hedges, C., et al. 2022, AJ, 163, 99
- Goffo et al. (2023) Goffo, E., Gandolfi, D., Egger, J. A., et al. 2023, ApJ, 955, L3
- Gronoff et al. (2020) Gronoff, G., Arras, P., Baraka, S., et al. 2020, Journal of Geophysical Research (Space Physics), 125, e27639
- Gu & Chen (2023) Gu, P.-G. & Chen, H. 2023, The Astrophysical Journal Letters, 953, L27
- Guillot (2010) Guillot, T. 2010, A&A, 520, A27
- Guo (2019) Guo, J. H. 2019, ApJ, 872, 99
- Gupta & Schlichting (2020) Gupta, A. & Schlichting, H. E. 2020, MNRAS, 493, 792
- Hakim et al. (2018) Hakim, K., Rivoldini, A., Van Hoolst, T., et al. 2018, Icarus, 313, 61
- Haldemann et al. (2020) Haldemann, J., Alibert, Y., Mordasini, C., & Benz, W. 2020, A&A, 643, A105
- Hamano et al. (2013) Hamano, K., Abe, Y., & Genda, H. 2013, Nature, 497, 607
- Hemley et al. (1992) Hemley, R. J., Stixrude, L., Fei, Y., & Mao, H. K. 1992, Geophysical Monograph Series, 67, 183
- Hirschmann et al. (2012) Hirschmann, M. M., Withers, A. C., Ardia, P., & Foley, N. T. 2012, Earth and Planetary Science Letters, 345, 38
- Ho et al. (2024) Ho, C. S. K., Rogers, J. G., Van Eylen, V., Owen, J. E., & Schlichting, H. E. 2024, MNRAS, 531, 3698
- Ho & Van Eylen (2023) Ho, C. S. K. & Van Eylen, V. 2023, MNRAS, 519, 4056
- Hollenbach et al. (1994) Hollenbach, D., Johnstone, D., Lizano, S., & Shu, F. 1994, ApJ, 428, 654
- Holmberg & Madhusudhan (2024) Holmberg, M. & Madhusudhan, N. 2024, A&A, 683, L2
- Howell et al. (2014) Howell, S. B., Sobeck, C., Haas, M., et al. 2014, Publications of the Astronomical Society of the Pacific, 126, 398
- Hu et al. (2015) Hu, R., Seager, S., & Yung, Y. L. 2015, ApJ, 807, 8
- Hunten et al. (1987) Hunten, D. M., Pepin, R. O., & Walker, J. C. G. 1987, Icarus, 69, 532
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Ida & Lin (2008) Ida, S. & Lin, D. N. C. 2008, ApJ, 673, 487
- Ikoma et al. (2018) Ikoma, M., Elkins-Tanton, L., Hamano, K., & Suckale, J. 2018, Space Science Reviews, 214, 76
- Inaba & Ikoma (2003) Inaba, S. & Ikoma, M. 2003, A&A, 410, 711
- Inaba et al. (2001) Inaba, S., Tanaka, H., Nakazawa, K., Wetherill, G. W., & Kokubo, E. 2001, Icarus, 149, 235
- Inamdar & Schlichting (2015) Inamdar, N. K. & Schlichting, H. E. 2015, MNRAS, 448, 1751
- Izidoro et al. (2022) Izidoro, A., Schlichting, H. E., Isella, A., et al. 2022, ApJ, 939, L19
- Jin & Mordasini (2018) Jin, S. & Mordasini, C. 2018, ApJ, 853, 163
- Jin et al. (2014) Jin, S., Mordasini, C., Parmentier, V., et al. 2014, ApJ, 795, 65
- Johnstone (2020) Johnstone, C. P. 2020, ApJ, 890, 79
- Katz et al. (2003) Katz, R. F., Spiegelman, M., & Langmuir, C. H. 2003, Geochemistry, Geophysics, Geosystems, 4, 1073
- Kempton & Knutson (2024) Kempton, E. M. R. & Knutson, H. A. 2024, Reviews in Mineralogy and Geochemistry, 90, 411
- Kempton et al. (2023) Kempton, E. M. R., Zhang, M., Bean, J. L., et al. 2023, Nature, 620, 67
- Kippenhahn et al. (2012) Kippenhahn, R., Weigert, A., & Weiss, A. 2012, Stellar Structure and Evolution (Berlin, Heidelberg: Springer)
- Kite et al. (2019) Kite, E. S., Fegley, Jr., B., Schaefer, L., & Ford, E. B. 2019, ApJ, 887, L33
- Kite et al. (2020) Kite, E. S., Fegley, Jr., B., Schaefer, L., & Ford, E. B. 2020, ApJ, 891, 111
- Kite & Schaefer (2021) Kite, E. S. & Schaefer, L. 2021, ApJ, 909, L22
- Komacek et al. (2019) Komacek, T. D., Showman, A. P., & Parmentier, V. 2019, ApJ, 881, 152
- Kopparapu et al. (2014) Kopparapu, R. K., Ramirez, R. M., SchottelKotte, J., et al. 2014, ApJ, 787, L29
- Kossakowski et al. (2021) Kossakowski, D., Kemmer, J., Bluhm, P., et al. 2021, A&A, 656, A124
- Kubyshkina et al. (2018) Kubyshkina, D., Fossati, L., Erkaev, N. V., et al. 2018, A&A, 619, A151
- Kubyshkina & Fossati (2021) Kubyshkina, D. I. & Fossati, L. 2021, Research Notes of the American Astronomical Society, 5, 74
- Kunimoto et al. (2022) Kunimoto, M., Winn, J., Ricker, G. R., & Vanderspek, R. K. 2022, AJ, 163, 290
- Lacedelli et al. (2024) Lacedelli, G., Pallè, E., Luque, R., et al. 2024, ArXiv e-prints
- Lammer et al. (2020) Lammer, H., Scherf, M., Kurokawa, H., et al. 2020, Space Science Reviews, 216, 74
- Leleu et al. (2024a) Leleu, A., Delisle, J.-B., Burn, R., et al. 2024a, A&A, 687, L1
- Leleu et al. (2024b) Leleu, A., Delisle, J. B., Delrez, L., et al. 2024b, A&A, 688, A211
- Lichtenberg (2021) Lichtenberg, T. 2021, ApJ, 914, L4
- Lichtenberg et al. (2019) Lichtenberg, T., Golabek, G. J., Burn, R., et al. 2019, Nature Astronomy, 3, 307
- Lichtenberg et al. (2023) Lichtenberg, T., Schaefer, L. K., Nakajima, M., & Fischer, R. A. 2023, in Protostars and Planets VII, Astronomical Society of the Pacific Conference Series, ed. Inutsuka, S. and Aikawa, Y. and Muto, T. and Tomida, K. and Tamura, M., Vol. 534, Kyoto, 907
- Lin et al. (1996) Lin, D. N. C., Bodenheimer, P., & Richardson, D. C. 1996, Nature, 380, 606
- Linder et al. (2019) Linder, E. F., Mordasini, C., Mollière, P., et al. 2019, A&A, 623, A85
- Lissauer et al. (2023) Lissauer, J. J., Batalha, N. M., & Borucki, W. J. 2023, in Protostars and Planets VII, Astronomical Society of the Pacific Conference Series, ed. Inutsuka, S. and Aikawa, Y. and Muto, T. and Tomida, K. and Tamura, M., Vol. 534, Kyoto, 839
- Lopez (2017) Lopez, E. D. 2017, MNRAS, 472, 245
- Lopez & Fortney (2013) Lopez, E. D. & Fortney, J. J. 2013, ApJ, 776, 2
- Loyd et al. (2020) Loyd, R. O. P., Shkolnik, E. L., Schneider, A. C., et al. 2020, ApJ, 890, 23
- Luo et al. (2024) Luo, H., Dorn, C., & Deng, J. 2024, Majority of Water Hides Deep in the Interiors of Exoplanets
- Luque et al. (2022) Luque, R., Fulton, B. J., Kunimoto, M., et al. 2022, A&A, 664, A199
- Luque et al. (2023) Luque, R., Osborn, H. P., Leleu, A., et al. 2023, Nature, 623, 932
- Luque & Pallé (2022) Luque, R. & Pallé, E. 2022, Science, 377, 1211
- MacDougall et al. (2023) MacDougall, M. G., Gilbert, G. J., & Petigura, E. A. 2023, AJ, 166, 61
- Makino et al. (1998) Makino, J., Fukushige, T., Funato, Y., & Kokubo, E. 1998, New Astronomy, 3, 411
- Malsky et al. (2023) Malsky, I., Rogers, L., Kempton, E. M.-R., & Marounina, N. 2023, Nature Astronomy, 7, 57
- Malsky & Rogers (2020) Malsky, I. & Rogers, L. A. 2020, ApJ, 896, 48
- Marboeuf et al. (2014a) Marboeuf, U., Thiabaud, A., Alibert, Y., Cabral, N., & Benz, W. 2014a, A&A, 570, A36
- Marboeuf et al. (2014b) Marboeuf, U., Thiabaud, A., Alibert, Y., Cabral, N., & Benz, W. 2014b, A&A, 570, A35
- Mason & Marrero (1970) Mason, E. A. & Marrero, T. R. 1970, in Advances in Atomic and Molecular Physics, ed. D. R. Bates & I. Esterrnan, Vol. 6 (Academic Press), 155–232
- McDonald et al. (2019) McDonald, G. D., Kreidberg, L., & Lopez, E. 2019, ApJ, 876, 22
- Miguel et al. (2020) Miguel, Y., Cridland, A., Ormel, C., Fortney, J., & Ida, S. 2020, MNRAS, 491, 1998
- Mishra et al. (2021) Mishra, L., Alibert, Y., Leleu, A., et al. 2021, A&A, 656, A74
- Moldenhauer et al. (2022) Moldenhauer, T. W., Kuiper, R., Kley, W., & Ormel, C. W. 2022, A&A, 661, A142
- Mordasini (2014) Mordasini, C. 2014, A&A, 572, A118
- Mordasini (2020) Mordasini, C. 2020, A&A, 638, A52
- Mordasini et al. (2009) Mordasini, C., Alibert, Y., & Benz, W. 2009, A&A, 501, 1139
- Mordasini et al. (2012a) Mordasini, C., Alibert, Y., Georgy, C., et al. 2012a, A&A, 547, A112
- Mordasini et al. (2012b) Mordasini, C., Alibert, Y., Klahr, H., & Henning, T. 2012b, A&A, 547, A111
- Mordasini & Burn (2024) Mordasini, C. & Burn, R. 2024, Reviews in Mineralogy and Geochemistry, 90, 55
- Mordasini et al. (2016) Mordasini, C., van Boekel, R., Mollière, P., Henning, Th., & Benneke, B. 2016, ApJ, 832, 41
- Mousis et al. (2020) Mousis, O., Deleuil, M., Aguichine, A., et al. 2020, ApJ, 896, L22
- Murray-Clay et al. (2009) Murray-Clay, R. A., Chiang, E. I., & Murray, N. 2009, ApJ, 693, 23
- Musella et al. (2019) Musella, R., Mazevet, S., & Guyot, F. 2019, Physical Review B, 99, 064110
- Nixon et al. (2024) Nixon, M. C., Piette, A. A. A., Kempton, E. M. R., et al. 2024, ApJ, 970, L28
- Odert et al. (2018) Odert, P., Lammer, H., Erkaev, N. V., et al. 2018, Icarus, 307, 327
- Orell-Miquel et al. (2023) Orell-Miquel, J., Nowak, G., Murgas, F., et al. 2023, A&A, 669, A40
- Osborn et al. (2021) Osborn, A., Armstrong, D. J., Cale, B., et al. 2021, MNRAS, 507, 2782
- Owen & Wu (2013) Owen, J. E. & Wu, Y. 2013, ApJ, 775, 105
- Paardekooper et al. (2023) Paardekooper, S., Dong, R., Duffell, P., et al. 2023, in ASP Conference Series, Vol. 534, Protostars and Planets VII, Astronomical Society of the Pacific Conference Series, ed. Inutsuka, S. and Aikawa, Y. and Muto, T. and Tomida, K. and Tamura, M., Kyoto, 685
- Paardekooper et al. (2011) Paardekooper, S.-J., Baruteau, C., & Kley, W. 2011, MNRAS, 410, 293
- Parc et al. (2024) Parc, L., Bouchy, F., Venturini, J., Dorn, C., & Helled, R. 2024, A&A, 688, A59
- Parviainen et al. (2024) Parviainen, H., Luque, R., & Palle, E. 2024, MNRAS, 527, 5693
- Passegger et al. (2024) Passegger, V. M., Suárez Mascareño, A., Allart, R., et al. 2024, A&A, 684, A22
- Pavlov (2019) Pavlov, A. V. 2019, Surveys in Geophysics, 40, 247
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011Journal of Machine Learning Research, 12, 2825
- Petigura et al. (2018) Petigura, E. A., Marcy, G. W., Winn, J. N., et al. 2018, AJ, 155, 89
- Petigura et al. (2022) Petigura, E. A., Rogers, J. G., Isaacson, H., et al. 2022, AJ, 163, 179
- Piaulet-Ghorayeb et al. (2024) Piaulet-Ghorayeb, C., Benneke, B., Radica, M., et al. 2024, JWST/NIRISS Reveals the Water-Rich ”Steam World” Atmosphere of GJ 9827 d
- Pierrehumbert (2023) Pierrehumbert, R. T. 2023, ApJ, 944, 20
- Richert et al. (2018) Richert, A. J. W., Getman, K. V., Feigelson, E. D., et al. 2018, MNRAS, 477, 5191
- Ricker et al. (2015) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2015, Journal of Astronomical Telescopes, Instruments, and Systems, 1, 014003
- Ricker et al. (2014) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2014, Journal of Astronomical Telescopes, Instruments, and Systems, 1, 014003
- Rogers et al. (2024) Rogers, J. G., Dorn, C., Raj, V. A., Schlichting, H. E., & Young, E. D. 2024, Most Super-Earths Have Less Than 3% Water
- Rogers & Owen (2021) Rogers, J. G. & Owen, J. E. 2021, MNRAS, 503, 1526
- Rogers et al. (2023) Rogers, J. G., Schlichting, H. E., & Owen, J. E. 2023, ApJ, 947, L19
- Sarkis et al. (2021) Sarkis, P., Mordasini, C., Henning, T., Marleau, G. D., & Mollière, P. 2021, A&A, 645, A79
- Schlecker et al. (2022) Schlecker, M., Burn, R., Sabotta, S., et al. 2022, A&A, 664, A180
- Schlichting & Young (2022) Schlichting, H. E. & Young, E. D. 2022, The Planetary Science Journal, 3, 127
- Schulik & Booth (2023) Schulik, M. & Booth, R. A. 2023, MNRAS, 523, 286
- Seager et al. (2007) Seager, S., Kuchner, M., Hier-Majumder, C. A., & Militzer, B. 2007, ApJ, 669, 1279
- Shah et al. (2021) Shah, O., Alibert, Y., Helled, R., & Mezger, K. 2021, A&A, 646, A162
- Silverberg et al. (2020) Silverberg, S. M., Wisniewski, J. P., Kuchner, M. J., et al. 2020, ApJ, 890, 106
- Sossi et al. (2023) Sossi, P. A., Tollan, P. M. E., Badro, J., & Bower, D. J. 2023, Earth and Planetary Science Letters, 601, 117894
- Stixrude (2014) Stixrude, L. 2014, Philosophical Transactions of the Royal Society of London Series A, 372, 20130076
- Stixrude & Lithgow-Bertelloni (2022) Stixrude, L. & Lithgow-Bertelloni, C. 2022, Geophysical Journal International, 228, 1119
- Thiabaud et al. (2014) Thiabaud, A., Marboeuf, U., Alibert, Y., et al. 2014, A&A, 562, A27
- Turbet et al. (2020) Turbet, M., Bolmont, E., Ehrenreich, D., et al. 2020, A&A, 638, A41
- Tychoniec et al. (2018) Tychoniec, Ł., Tobin, J. J., Karska, A., et al. 2018, The Astrophysical Journal Supplement Series, 238, 19
- Ulmer-Moll et al. (2023) Ulmer-Moll, S., Osborn, H. P., Tuson, A., et al. 2023, A&A, 674, A43
- Van Eylen et al. (2018) Van Eylen, V., Agentoft, C., Lundkvist, M. S., et al. 2018, MNRAS, 479, 4786
- Van Eylen et al. (2021) Van Eylen, V., Astudillo-Defru, N., Bonfils, X., et al. 2021, MNRAS, 507, 2154
- Vazan et al. (2022) Vazan, A., Sari, R., & Kessel, R. 2022, ApJ, 926, 150
- Venturini et al. (2020) Venturini, J., Guilera, O. M., Haldemann, J., Ronco, M. P., & Mordasini, C. 2020, A&A, 643, L1
- Venturini et al. (2024) Venturini, J., Ronco, M. P., Guilera, O. M., et al. 2024, A&A, 686, L9
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wallack et al. (2024) Wallack, N. L., Batalha, N. E., Alderson, L., et al. 2024, JWST COMPASS: A NIRSpec/G395H Transmission Spectrum of the Sub-Neptune TOI-836c
- Wang et al. (2023) Wang, Y., Ormel, C. W., Huang, P., & Kuiper, R. 2023, MNRAS, 523, 6186
- Ward (1997) Ward, W. R. 1997, Icarus, 126, 261
- Yoshida et al. (2022) Yoshida, T., Terada, N., Ikoma, M., & Kuramoto, K. 2022, ApJ, 934, 137
- Zahnle et al. (1990) Zahnle, K., Kasting, J. F., & Pollack, J. B. 1990, Icarus, 84, 502
- Zahnle & Kasting (1986) Zahnle, K. J. & Kasting, J. F. 1986, Icarus, 68, 462
- Zeng et al. (2019) Zeng, L., Jacobsen, S. B., Sasselov, D. D., et al. 2019, Proceedings of the National Academy of Sciences, 116, 9723
- Zeng et al. (2016) Zeng, L., Sasselov, D. D., & Jacobsen, S. B. 2016, ApJ, 819, 127
Appendix A An analytic model for fractionation
For large escape rates, the analytic estimate by Hunten et al. (1987) should apply (Zahnle et al. 1990). They derived a fractionation factor
where is the mass of the planet, is the temperature in the atmosphere assumed in this model to be isothermal, is the binary diffusion coefficient (Mason & Marrero 1970; Zahnle & Kasting 1986; Pavlov 2019), is the effective planetary radius, and (respectively ) is the escaping flux of hydrogen (oxygen) particles per surface such that the total particle loss ().
@@ Appendix A An analytic model for fractionation
For the optimistic fractionation model, we determine the base of the flow analogous to literature on disk photoevaporation (Hollenbach et al. 1994) by requiring that the mass flux launched as a decoupled wind at the absorption layer of high-energy radiation (Murray-Clay et al. 2009) is sustained from a deeper ’base’ radius of the flux. This can be written as
where , with for mass fractions of hydrogen , helium and heavy elements assumed to be water . This averaging assumes that particles are singly ionized. For example H2O is present as H+, O+, and 3 electrons with a total weight of 18 u distributed to 6 particles, thus an average weight of 3 u per particle results. The sound speed is an order-of-magnitude estimate of average velocity with which the ionized particles in the wind are launched. The base radius can be located outside of the resolved structure of the envelope which sometimes requires extrapolation. Then, the number density ratio is estimated assuming individual hydrostatic distributions for O and H with scale heights . This implies no interaction between O and H, thus a lower limit for , respectively an upper limit for the fractionation effect. More detailed considerations should include different species (Odert et al. 2018), the effect of turbulent mixing (e.g. Charnay et al. 2015; Komacek et al. 2019), ionization and the resulting interactions with ions (Hu et al. 2015; Guo 2019), and the deviations of detailed numerical studies from the analytical approach (Johnstone 2020; Schulik & Booth 2023).

### v2

@@ Acknowledgements.
We thank J. Venturini, C. Mordasini, Th. Henning, H. Klahr, and E. Pallé for fruitful discussion. We further thank the contributions and discussions of all participants of the ”Density Matters” Ringberg workshop which helped to shape this work. Finally, we thank the anonymous referee for insightful comments which helped to improve the quality of analysis and manuscript. R.B. acknowledges the financial support from DFG under Germany’s Excellence Strategy EXC 2181/1-390900948, Exploratory project EP 8.4 (the Heidelberg STRUCTURES Excellence Cluster) and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under grant agreement No. 832428-Origins and via the ERC Advanced Grant “TiPPi – Turbulence, Pebbles and Planetesimals: The Origin of Minor Bodies in the Solar System” (project ID 855130, PI: Klahr). R.L. acknowledges support provided by NASA through the NASA Hubble Fellowship grant #HST-HF2-51559.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. C.D. acknowledges support from the Swiss National Science Foundation under grant TMSGI2_211313. Parts of this work has been carried out within the framework of the NCCR PlanetS supported by the Swiss National Science Foundation under grants 51NF40_182901 and 51NF40_205606. Calculation were performed on the Horus cluster at the University of Bern and on the computation facilities at the Max Planck Institute for Astronomy. The plots shown in this work were generated using matplotlib (Hunter2007) and the analysis made use of the scikit-learn (scikit-learn) and scipy (scipy) packages.Appendix A An analytic model for fractionation
For large escape rates, the analytic estimate by Hunten1987 should apply (Zahnle1990). They derived a fractionation factor
where is the mass of the planet, is the temperature in the atmosphere assumed in this model to be isothermal, is the binary diffusion coefficient (Mason1970; Zahnle1986; Pavlov2019), is the effective planetary radius, and (respectively ) is the escaping flux of hydrogen (oxygen) particles per surface such that the total particle loss ().
@@
For the optimistic fractionation model, we determine the base of the flow analogous to literature on disk photoevaporation (Hollenbach1994) by requiring that the mass flux launched as a decoupled wind at the absorption layer of high-energy radiation (Murray-Clay2009) is sustained from a deeper ’base’ radius of the flux. This can be written as
where , with for mass fractions of hydrogen , helium and heavy elements assumed to be water . This averaging assumes that particles are singly ionized. For example H2O is present as H+, O+, and 3 electrons with a total weight of 18 u distributed to 6 particles, thus an average weight of 3 u per particle results. The sound speed is an order-of-magnitude estimate of average velocity with which the ionized particles in the wind are launched. The base radius can be located outside of the resolved structure of the envelope which sometimes requires extrapolation. Then, the number density ratio is estimated assuming individual hydrostatic distributions for O and H with scale heights . This implies no interaction between O and H, thus a lower limit for , respectively an upper limit for the fractionation effect. More detailed considerations should include different species (Odert2018), the effect of turbulent mixing (e.g. Charnay2015; Komacek2019), ionization and the resulting interactions with ions (Hu2015; Guo2019), and the deviations of detailed numerical studies from the analytical approach (Johnstone2020; Schulik2023).